# AI-driven Conservative-to-Primitive Conversion in Hybrid Piecewise Polytropic and Tabulated Equations of State

### v1

@@ Abstract
We present a novel AI-based approach to accelerate conservative-to-primitive inversion in relativistic hydrodynamics simulations, focusing on hybrid piecewise polytropic and tabulated equations of state. Traditional root-finding methods are computationally intensive, particularly in large-scale simulations. To address this, we employ feedforward neural networks (NNC2PS and NNC2PL), trained in PyTorch and optimized for GPU inference using NVIDIA TensorRT, achieving significant speedups with minimal loss in accuracy. The NNC2PS model achieves $L_{1}$ and $L_{\infty}$ errors of $4.54\times 10^{-7}$ and $3.44\times 10^{-6}$, respectively, with the NNC2PL model yielding even lower error values. TensorRT optimization ensures high accuracy, with FP16 quantization offering 7x faster performance than traditional root-finding methods. Our AI models outperform conventional CPU solvers, demonstrating enhanced inference times, particularly for large datasets. We release the scientific software developed for this work, enabling the validation and extension of our findings. These results highlight the potential of AI, combined with GPU optimization, to significantly improve the efficiency and scalability of numerical relativity simulations.
preprint: APS/123-QED
I Introduction
@@
In numerical relativity, accurately modeling astrophysical systems such as neutron star mergers relies on solving the equations of relativistic hydrodynamics, which involve the inversion of conservative-to-primitive (C2P) variable relations [1, 2, 3]. This process, central to simulations of general relativistic hydrodynamics (GRHD), typically requires computationally expensive root-finding algorithms, such as Newton-Raphson methods, and interpolation of complex, multi-dimensional equations of state (EOS) tables [4, 5]. These methods, while robust, incur significant computational costs and can lead to inefficiencies, particularly in large-scale simulations, where up to billions of C2P calls may be required per time step. This limitation is especially pronounced in simulations that aim to produce accurate gravitational waveforms for sources that may be detected by gravitational wave detectors like Advanced LIGO, where the accuracy of the results must be achieved efficiently and within feasible computational times [6, 7, 8, 9, 10, 11, 12, 13].
@@
Recent advances have explored the potential of machine learning (ML) to replace traditional root-finding approaches in this context [14]. Specifically, neural networks have shown promise in accelerating the C2P inversion process while maintaining high accuracy [14]. Building on this, the present work introduces a novel approach that leverages ML to accelerate the recovery of primitive variables from conserved variables in relativistic hydrodynamics simulations, with particular focus on hybrid piecewise polytropic and tabulated EOS. These EOS models provide more realistic descriptions of the dense interior of neutron stars, yet their complexity makes the traditional C2P procedure very computationally expensive.
@@
We present a suite of feedforward neural networks trained to directly map conserved variables to primitive variables, bypassing the need for traditional iterative solvers. In particular, we employ a hybrid approach, utilizing the flexibility of neural networks to handle the challenges posed by complex EOS models. Our models are implemented using modern deep learning tools, such as PyTorch, and optimized for GPU inference with NVIDIA TensorRT  [15, 16]. Through comprehensive performance benchmarking, we demonstrate that our approach significantly outperforms traditional numerical methods in terms of speed, particularly when using mixed-precision deployment on modern hardware accelerators like the NVIDIA A100 GPU.
We also quantify the scalability of our ML models, demonstrating near-linear scaling for large datasets and substantial reductions in inference times, even for datasets containing hundreds of thousands of data points. These results highlight the potential of ML-based methods to enable high-accuracy, high-throughput simulations, positioning them as an essential tool for next-generation numerical relativity codes. By accelerating the C2P inversion, our approach facilitates the more efficient generation of gravitational waveforms, directly contributing to the ongoing efforts to better understand and characterize the behavior of dense astrophysical systems. Accurate numerical relativity simulations of binary neutron star mergers are critical for advancing our understanding of the complex dynamics of these events. These simulations provide unique insights into the highly nonlinear interplay between gravitational waves, matter, and nuclear physics, particularly in the extreme environments surrounding such mergers [17, 18]. By comparing simulation results with gravitational wave observations, we can probe the EOS of dense nuclear matter, offering constraints on the properties of neutron star interiors [19, 20]. Moreover, these simulations enable the exploration of fundamental microphysics, such as the behavior of matter at supra-nuclear densities, shedding light on phenomena like hyperons, quark matter, and nuclear phase transitions [21, 22]. Thus, by enhancing computational efficiency, our approach not only accelerates waveform generation but also advances the ability to explore the intricate physics of compact binary mergers, enabling more comprehensive studies of these astrophysical phenomena.
@@
This article is structured as follows. Section II introduces the EOS considered in this study, along with the methodologies employed for designing, training, validating, and testing the ML models. In Section III, we present our key results, including an assessment of the accuracy of the ML models across different model types and quantization schemes. Additionally, we provide a comparison of the computational performance of the ML models relative to traditional root-finding methods. Finally, Section IV offers a summary of the findings and outlines potential avenues for future research.
II Methods
@@
We present a ML-based model with the potential to accelerate recovery of primitive variables from conserved variables in general relativistic hydrodynamics simulations, specifically focusing on scenarios employing hybrid piecewise polytropic EOS, and tabulated EOS. As in traditional approaches, this conversion requires inverting the conservative-to-primitive map, a process often reliant on computationally expensive root-finding algorithms. While previous work has demonstrated the success of machine learning for this task with the $\Gamma-$law EOS [23], here we investigate its application to hybrid piecewise polytropic EOS, which offers a more realistic representation of neutron star interiors, as well as the tabulated EOS, which incorporate current nuclear physics model of neutron matter. To evaluate the performance of our neural network, we use a traditional CPU-based root-finding algorithm (provided by the RePrimAnd library [24]) as a baseline for comparison. Our aim is to demonstrate the speed advantages of the neural network approach for conservative-to-primitive variable conversion. Our network is implemented using PyTorch, and the inference speed tests are performed using libtorch and NVIDIA TensorRT’s C++ API.
@@
$\mathbf{F}^{i}$ $=\left(D\left(v^{i}-\frac{\beta^{i}}{\alpha}\right),S_{j}\left(v^{i}-\frac{\beta^{i}}{\alpha}\right)+p\delta^{i}_{j},\right.$
$\quad\left.\tau\left(v^{i}-\frac{\beta^{i}}{\alpha}\right)+pv^{i}\right)\,,$
@@
We will first focus on the hybrid piecewise polytropic EOS. The hybrid piecewise polytropic EOS was introduced for simplified simulations of stellar collapse to model the stiffening of the nuclear EOS at nuclear density, and include thermal pressure during the postbounce phase [28]. In gravitational-wave science, it is more commonly used as described in Read et al. [29], where it enables gravitational-wave parameter estimation and waveform modeling by effectively capturing macroscopic neutron star observables with minimal parameters. The structure of this EOS consists of multiple cold polytropes, defined by parameters $K_{0},K_{1},\cdots,K_{\text{nsegments}-1}$ and $\Gamma_{0},\Gamma_{1},\cdots,\Gamma_{\text{nsegments}-1}$, where nsegments denotes the total number of segments. Additionally, it includes a thermal $\Gamma-$law component characterized by $\Gamma_{\text{th}}$. Continuity of pressure and internal energy across segments, in accordance with the first law of thermodynamics, is ensured after appropriately setting initial values for the polytropic indices, density breakpoints (denoted $\rho_{\text{breaks}})$, and other relevant parameters. In this context, pressure and specific internal energy components in each density interval are given by
$p_{\text{cold}}$ $=K_{i}\rho^{\Gamma_{i}},$
$\epsilon_{\text{cold}}$ $=a_{i}+\frac{K_{i}}{\Gamma_{i}-1}\rho^{\Gamma_{i}-1},$ (3)
$p_{\text{th}}$ $=(\Gamma_{\text{th}}-1)\rho(\epsilon-\epsilon_{\text{cold}}),$
$p$ $=p_{\text{th}}+p_{\text{cold}},$
where $a_{i}$ is segment-specific constant, and the rest mass density, $\rho$, is assumed to fall into the segments specified by each of $\rho_{\text{breaks}}$.
@@
II.1 Data
II.1.1 Piecewise Polytropic EOS Based Model Data
@@
We generate a dataset of 500,000 samples using geometrized units where $G=c=M_{\odot}=1$. Without loss of generality, we furthermore use a Minkoswki metric $g_{\mu\nu}=\operatorname{diag}(-1,+1,+1,+1)$. The rest-mass density, $\rho$, is sampled uniformly from $[2\times 10^{-5},2\times 10^{-3}]$, and the fluid’s three-velocity is assumed one-dimensional along the $x$-axis, sampled uniformly from $v_{x}\in[0,0.721)$. A four-segment piecewise polytrope is used, with segment-wise polytropic indices $\Gamma=[1.3569,3.0050,2.9880,2.8510]$. The first segment’s polytropic constant, $K_{0}$, is set to $8.9493\times 10^{-2}$. Subsequent polytropic constants, $K_{i}$, are determined by enforcing pressure continuity. Similarly, the first segment’s constant, $a_{0}$, is set to zero, while subsequent $a_{i}$ values ensure continuity of internal energy. The density breaks for the segments are specified at $\rho=2.3674\times 10^{-4}$, $8.1147\times 10^{-4}$, and $1.6191\times 10^{-3}$. The thermal component has an adiabatic index of $\Gamma_{\text{th}}=5/3$. Additionally, the thermal component of the specific internal energy, $\epsilon_{\text{th}}$, is sampled uniformly from $[0,2]$. A structured dataset is then constructed by converting the primitive variables to conserved variables using the standard relativistic hydrodynamic relations given in Eq. II. In this dataset, conserved variables serve as input features, and the pressure is the target variable. The resulting dataset is then split into training, validation, and test sets, with each set fully standardized to zero mean and unit variance to ensure equal contribution of all features during neural network training (Fig. 1).
II.1.2 Tabulated EOS Based Model Data
@@
To generate the training data for the tabulated EOS-based model, we sample from a provided EOS table and follow a procedure similar to the one described in Section II.1.1. We begin by reading in the EOS table, which contains the variables electron fraction ($Y_{e}$), temperature ($T$), rest-mass density ($\rho$), specific internal energy ($\epsilon$), and pressure ($p$). These quantities are stored in logarithmic form in the table and are extracted accordingly. For each data point, a random one-dimensioal three-velocity, $v_{x}$, is sampled from the interval $[0,0.721)$, and values are randomly chosen for electron fraction, temperature, and density from among the tabulated values. Using these, the corresponding values of $\rho$, $\epsilon$, and $p$ are then fetched from the EOS table. The primitive variables are then converted into conserved variables using standard relativistic hydrodynamics relations given in Eq. II. A total of 1,000,000 datapoints are generated using this process  [31]. Similar to the hybrid piecewise polytropic EOS-based model, the data is split into training, validation, and test sets, with each set fully standardized to zero mean and unit variance before being used for neural network training.
@@
II.2 Model architecture
II.2.1 Piecewise Polytropic EOS Based Model
For the hybrid piecewise polytropic EOS-based model, we tested two feedforward neural networks of varying complexity to represent the conservative-to-primitive variable transformation. Each network takes as input the three conserved variables $(D,S_{x},\tau)$ (Eq.II) and outputs the pressure $p$ (Eq.II), assuming the remaining momentum density components are zero for simplicity. After experimenting with multiple multi-layer perceptron (MLP) architectures, we identified two models that offered a good balance between accuracy, speed, and trainability. The smaller model, NNC2PS, features two hidden layers with 600 and 200 neurons, while the larger model, NNC2PL, contains five hidden layers with 1024, 512, 256, 128, and 64 neurons (Fig 2).
II.2.2 Tabulated EOS Based Model
For the tabulated EOS-based model, we use a single feedforward neural network, NNC2P_Tabulated, to achieve an inherently equivalent task with minor differences. This model takes as input the log-scaled variables $(\log D,\log S_{x},\log\tau,\log Y_{e})$ and outputs the log-scaled pressure, $\log p$ (Eq.II), assuming $S_{y}$ and $S_{z}$ are zero for simplicity as before. Using log-scaled inputs and outputs aligns with the format of the tabulated EOS values, which are also stored in logarithmic form to accommodate the typically large values of these physical quantities. This approach reduces the range of feature magnitudes, facilitating more stable learning dynamics and better alignment with the source data.
We explored several MLP architectures, varying in parameters, layers, and training strategies, to identify an optimal design for our task. Among these, a structure resembling NNC2PL, detailed in Section II.2.1 above, emerged as a robust choice. This architecture effectively balanced capacity and efficiency, enabling accurate learning of log-scaled pressure from tabulated EOS data (Fig. 2).
II.3 Training approach

### v2

# Machine learning-driven conservative-to-primitive conversion in hybrid piecewise polytropic and tabulated equations of state

@@ Abstract
We present a novel machine learning (ML) method to accelerate conservative-to-primitive inversion, focusing on hybrid piecewise polytropic and tabulated equations of state. Traditional root-finding techniques are computationally expensive, particularly for large-scale relativistic hydrodynamics simulations. To address this, we employ feedforward neural networks (NNC2PS and NNC2PL), trained in PyTorch and optimized for GPU inference using NVIDIA TensorRT, achieving significant speedups with minimal accuracy loss. The NNC2PS model achieves $L_{1}$ and $L_{\infty}$ errors of $4.54\times 10^{-7}$ and $3.44\times 10^{-6}$, respectively, while the NNC2PL model exhibits even lower error values. TensorRT optimization with mixed-precision deployment substantially accelerates performance compared to traditional root-finding methods. Specifically, the mixed-precision TensorRT engine for NNC2PS achieves inference speeds approximately 400 times faster than a traditional single-threaded CPU implementation for a dataset size of 1,000,000 points. Ideal parallelization across an entire compute node in the Delta supercomputer (Dual AMD 64 core 2.45 GHz Milan processors; and 8 NVIDIA A100 GPUs with 40 GB HBM2 RAM and NVLink) predicts a 25-fold speedup for TensorRT over an optimally-parallelized numerical method when processing 8 million data points. Moreover, the ML method exhibits sub-linear scaling with increasing dataset sizes. We release the scientific software developed, enabling further validation and extension of our findings. This work underscores the potential of ML, combined with GPU optimization and model quantization, to accelerate conservative-to-primitive inversion in relativistic hydrodynamics simulations.
1 Introduction
@@
In numerical relativity, accurately modeling astrophysical systems such as neutron star mergers [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14] relies on solving the equations of relativistic hydrodynamics, which involve the inversion of conservative-to-primitive (C2P) variable relations [15, 16, 17]. This process typically requires computationally expensive root-finding algorithms, such as Newton-Raphson methods, and interpolation of complex, multi-dimensional equations of state (EOS) tables [18, 19]. These methods, while robust, incur significant computational costs and can lead to inefficiencies, particularly in large-scale simulations, where up to billions of C2P calls may be required per time step.
@@
In view of these considerations, and taking into account the advent of GPU-based exascale supercomputers such as Aurora and Frontier, and ongoing efforts to port relativistic hydrodynamics software into GPUs [20, 21, 22], this work explores the use of machine learning (ML) algorithms that leverage GPU-accelerated computing for C2P conversion. CPU based algorithms for C2P conversion typically involve an iterative non-linear root finder, for which the number of iterations required to achieve a given target accuracy depends on the input data, resulting in different runtimes for different points of numerical grid. This limits the potential to use SIMD (for CPUs) or SIMT (for GPUs) parallelism, reducing the effective rate of conversion achievable using these schemes. A ML approach with its more predictable runtime and regular memory access pattern may help alleviate these issues. Indeed, this work is motivated by recent studies that have explored the potential of ML to replace traditional root-finding approaches for C2P inversion [23]. Specifically, neural networks have shown promise in accelerating the C2P inversion process while maintaining high accuracy [23]. Building on this, the present work introduces a novel approach that leverages ML to accelerate the recovery of primitive variables from conserved variables in relativistic hydrodynamics simulations, with particular focus on hybrid piecewise polytropic and tabulated EOS. These EOS models provide more realistic descriptions of the dense interior of neutron stars, yet their complexity makes the traditional C2P procedure very computationally expensive.
@@
To help address these computational challenges, we present a suite of feedforward neural networks trained to directly map conserved variables to primitive variables, bypassing the need for traditional iterative solvers. In particular, we employ a hybrid approach, utilizing the flexibility of neural networks to handle the challenges posed by complex EOS models. Our models are implemented using modern deep learning tools, such as PyTorch, and optimized for GPU inference with NVIDIA TensorRT  [24, 25]. Through comprehensive performance benchmarking, we demonstrate that our approach significantly outperforms traditional numerical methods in terms of speed, particularly when using mixed-precision deployment on modern hardware accelerators like NVIDIA A100 GPUs in the Delta supercomputer.
We evaluate the scalability of our ML models by comparing their inference performance against a single-threaded CPU implementation of a traditional numerical method from the RePrimAnd library [26]. The benchmark was conducted on a Delta supercomputer compute node, featuring dual AMD 64-core 2.45 GHz Milan processors, 8 NVIDIA A100 GPUs (40 GB HBM2 RAM), and NVLink. For dataset sizes ranging from 25,000 to 1,000,000 points, the numerical method exhibited linear scaling of inference time. In contrast, TensorRT-optimized and TorchScript-based neural networks achieved substantially faster inference, typically demonstrating sub-linear scaling. We investigate two feedforward neural network architectures: a smaller network (NNC2PS) and a larger one (NNC2PL). Notably, mixed-precision TensorRT engines delivered impressive performance, with the NNC2PS engine processing 1,000,000 points in 8.54 ms, compared to 3,490 ms for the numerical method. Ideal parallelization across the entire node (64 CPU cores that support up to 128 threads, and 8 GPUs) suggests a 25-fold speedup for TensorRT over the optimally parallelized numerical method when processing 8 million points. These results demonstrate the scalability and efficiency of our ML-based methods, offering significant improvements for high-throughput numerical relativistic hydrodynamics simulations.
@@
This article is structured as follows. Section 2 introduces the EOS considered in this study, along with the methodologies employed for designing, training, validating, and testing the ML models. In Section 3, we present our key results, including an assessment of the accuracy of the ML models across different model types and quantization schemes. Additionally, we provide a comparison of the computational performance of the ML models relative to traditional root-finding methods. Finally, Section 4 offers a summary of the findings and outlines potential avenues for future research.
2 Methods
@@
We present a ML-based model with the potential to accelerate recovery of primitive variables from conserved variables in general relativistic hydrodynamics (GRHD) simulations, specifically focusing on scenarios employing hybrid piecewise polytropic EOS, and tabulated EOS. As in traditional approaches, this conversion requires inverting the conservative-to-primitive map, a process often reliant on computationally expensive root-finding algorithms. While previous work has demonstrated the success of machine learning for this task with the $\Gamma-$law EOS [23], here we investigate its application to hybrid piecewise polytropic EOS, which offers a more realistic representation of neutron star interiors, as well as the tabulated EOS, which incorporate current nuclear physics model of neutron matter. To evaluate the performance of our neural network, we use a traditional CPU-based root-finding algorithm (provided by the RePrimAnd library) as a baseline for comparison. Our aim is to demonstrate the speed advantages of the neural network approach for conservative-to-primitive variable conversion. Our network is implemented using PyTorch, and the inference speed tests are performed using libtorch and NVIDIA TensorRT’s C++ API.
@@
$\mathbf{F}^{i}=\left(D\left(v^{i}-\frac{\beta^{i}}{\alpha}\right),\;S_{j}\left(v^{i}-\frac{\beta^{i}}{\alpha}\right)+p\delta^{i}_{j},\;\right.\\ \left.\tau\left(v^{i}-\frac{\beta^{i}}{\alpha}\right)+pv^{i}\right)\,,$ (2)
@@
We will first focus on the hybrid piecewise polytropic EOS. The hybrid piecewise polytropic EOS was introduced for simplified simulations of stellar collapse to model the stiffening of the nuclear EOS at nuclear density, and include thermal pressure during the postbounce phase [30]. In gravitational-wave science, it is more commonly used as described in Read et al. [31], where it enables gravitational-wave parameter estimation and waveform modeling by effectively capturing macroscopic neutron star observables with minimal parameters. The structure of this EOS consists of multiple cold polytropes, defined by parameters $K_{0},K_{1},\cdots,K_{\mathrm{segments}-1}$ and $\Gamma_{0},\Gamma_{1},\cdots,\Gamma_{\mathrm{nsegments}-1}$, where nsegments denotes the total number of segments. Additionally, it includes a thermal $\Gamma-$law component characterized by $\Gamma_{\mathrm{th}}$. Continuity of pressure and internal energy across segments, in accordance with the first law of thermodynamics, is ensured after appropriately setting initial values for the polytropic indices, density breakpoints (denoted $\rho_{\mathrm{breaks}})$, and other relevant parameters. In this context, pressure and specific internal energy components in each density interval are given by
$p_{\mathrm{cold}}$ $=K_{i}\rho^{\Gamma_{i}},$
$\epsilon_{\mathrm{cold}}$ $=a_{i}+\frac{K_{i}}{\Gamma_{i}-1}\rho^{\Gamma_{i}-1},$ (4)
$p_{\mathrm{th}}$ $=(\Gamma_{\mathrm{th}}-1)\rho(\epsilon-\epsilon_{\mathrm{cold}}),$
$p$ $=p_{\mathrm{th}}+p_{\mathrm{cold}},$
where $a_{i}$ is segment-specific constant, and the rest mass density, $\rho$, is assumed to fall into the segments specified by each of $\rho_{\mathrm{breaks}}$.
@@
2.1 Data
2.1.1 Piecewise Polytropic EOS Based Model Data
@@
We generate a dataset of 500,000 samples using geometrized units where $G=c=M_{\odot}=1$. Without loss of generality, we furthermore use a Minkoswki metric $g_{\mu\nu}=\mathrm{diag}(-1,+1,+1,+1)$. The rest-mass density, $\rho$, is sampled uniformly from $[2\times 10^{-5},2\times 10^{-3}]$, and the fluid’s three-velocity is assumed one-dimensional along the $x$-axis, sampled uniformly from $v_{x}\in(0,0.721)$. Following Ref. [31], we use an SLy four-segment piecewise polytropic EOS with segment-wise polytropic indices $\Gamma=[1.3569,3.0050,2.9880,2.8510]$. The first segment’s polytropic constant, $K_{0}$, is set to $8.9493\times 10^{-2}$. Subsequent polytropic constants, $K_{i}$, are determined by enforcing pressure continuity. Similarly, the first segment’s constant, $a_{0}$, is set to zero, while subsequent $a_{i}$ values ensure continuity of internal energy. The density breaks for the segments are specified at $\rho=2.3674\times 10^{-4}$, $8.1147\times 10^{-4}$, and $1.6191\times 10^{-3}$. The thermal component has an adiabatic index of $\Gamma_{\mathrm{th}}=5/3$. Additionally, the thermal component of the specific internal energy, $\epsilon_{\mathrm{th}}$, is sampled uniformly from $[0,2]$. A structured dataset is then constructed by converting the primitive variables to conserved variables using the standard relativistic hydrodynamic relations given in Eq. 2. In this dataset, conserved variables serve as input features, and the pressure is the target variable. The resulting dataset is then split into training, validation, and test sets, with each set fully standardized to zero mean and unit variance to ensure equal contribution of all features during neural network training (Fig. 1).
2.1.2 Tabulated EOS Based Model Data
@@
To generate the training data for the tabulated EOS-based model, we sample from a provided EOS table and follow a procedure similar to the one described in Section 2.1.1. We begin by reading in the EOS table, which contains the variables electron fraction ($Y_{e}$), temperature ($T$), rest-mass density ($\rho$), specific internal energy ($\epsilon$), and pressure ($p$). These quantities are stored in logarithmic form in the table and are extracted accordingly. For each data point, a random one-dimensioal three-velocity, $v_{x}$, is sampled from the interval $(0,0.721)$, and values are randomly chosen for electron fraction, temperature, and density from among the tabulated values. Using these, the corresponding values of $\rho$, $\epsilon$, and $p$ are then fetched from the EOS table. The primitive variables are then converted into conserved variables using standard relativistic hydrodynamics relations given in Eq. 2. A total of 1,000,000 datapoints are generated using this process  [33]. Similar to the hybrid piecewise polytropic EOS-based model, the data is split into training, validation, and test sets, with each set fully standardized to zero mean and unit variance before being used for neural network training.
2.2 Model architecture
2.2.1 Piecewise Polytropic EOS Based Model
For the hybrid piecewise polytropic EOS-based model, we tested two feedforward neural networks of varying complexity to represent the conservative-to-primitive variable transformation. Each network takes as input the three conserved variables $(D,S_{x},\tau)$ (Eq. 2) and outputs the pressure $p$ (Eq. 2), assuming the remaining momentum density components are zero for simplicity. After experimenting with multiple multi-layer perceptron (MLP) architectures, we identified two models that offered a good balance between accuracy, speed, and trainability. The smaller model, NNC2PS, features two hidden layers with 600 and 200 neurons, while the larger model, NNC2PL, contains five hidden layers with 1024, 512, 256, 128, and 64 neurons (Fig 2).
2.2.2 Tabulated EOS Based Model
For the tabulated EOS-based model, we use a single feedforward neural network, NNC2P_Tabulated, to achieve an inherently equivalent task with minor differences. This model takes as input the log-scaled variables $(\log D,\log S_{x},\log\tau,\log Y_{e})$ and outputs the log-scaled pressure, $\log p$ (Eq. 2), assuming $S_{y}$ and $S_{z}$ are zero for simplicity as before. Using log-scaled inputs and outputs aligns with the format of the tabulated EOS values, which are also stored in logarithmic form to accommodate the typically large values of these physical quantities. This approach reduces the range of feature magnitudes, facilitating more stable learning dynamics and better alignment with the source data.
We explored several MLP architectures, varying in parameters, layers, and training strategies, to identify an optimal design for our task. Among these, a structure resembling NNC2PL, detailed in Section 2.2.1 above, emerged as a robust choice. This architecture effectively balanced capacity and efficiency, enabling accurate learning of log-scaled pressure from tabulated EOS data (Fig. 2).
@@ 2.2.2 Tabulated EOS Based Model
2.3 Training approach